OpenStudy (anonymous):

how to factor differences of squares? a2 - 4

OpenStudy (bahrom7893):

a^2 - 4 = a^2 - 2^2 = (a-2)(a+2)

OpenStudy (bahrom7893):

please click on Become a fan if I helped, thanks!

OpenStudy (anonymous):

What do you mean? Please fan me too. I'LL FAN BACK! :D

OpenStudy (anonymous):

factor each polynomial if possible. If the polynomial cannot be factored, write prime.